# CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction

### v1

@@ Abstract
The paper focuses on improving the interpretability of Grammatical Error Correction (GEC) metrics, which receives little attention in previous studies. To bridge the gap, we propose CLEME2.0, a reference-based evaluation strategy that can describe four elementary dimensions of GEC systems, namely hit-correction, error-correction, under-correction, and over-correction. They collectively contribute to revealing the critical characteristics and locating drawbacks of GEC systems. Evaluating systems by Combining these dimensions leads to high human consistency over other reference-based and reference-less metrics. Extensive experiments on 2 human judgement datasets and 6 reference datasets demonstrate the effectiveness and robustness of our method.111All the codes will be released after the peer review.
CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction
Jingheng Ye1††thanks: indicates equal contribution., Zishan Xu1∗, Yinghui Li1, Xuxin Cheng2, Linlin Song3, Qingyu Zhou, Hai-Tao Zheng1,4††thanks: Corresponding author: Hai-Tao Zheng. (E-mail: zheng.haitao@sz.tsinghua.edu.cn), Ying Shen5, Xin Su6 1Tsinghua University, 2Peking University 3Huazhong University of Science and Technology 4Peng Cheng Laboratory, 5Sun-Yat Sen University, 6Tencent {yejh22,xzs23}@mails.tsinghua.edu.cn
@@ 1 Introduction
Grammatical Error Correction (GEC) is the task of automatically detecting and correcting all grammatical errors in a given text Li et al. (2022a, b); Ma et al. (2022); Ye et al. (2022). A core component of any NLP tasks is the development of automatic metrics that can objectively measure model performance Bryant et al. (2023). However, proposing appropriate evaluation of GEC has long been a challenging task Madnani et al. (2011); Li et al. (2023a), due to the subjectivity Bryant and Ng (2015), complexity Mita et al. (2019) and subtlety Choshen and Abend (2018) of GEC Napoles et al. (2015).
Recent studies have been trying to develop GEC metrics that can achieve high correlations with human judgements Yoshimura et al. (2020a), with less attention paid to the interpretability of the automatic metrics. We define the interpretability of metrics as their ability to reveal the concerned characteristics of systems, which is vital in locating the drawbacks of a certain system. It is well-acknowledged that excellent GEC systems, which usually conform to the principle of minimal editing, should adhere to two gold principles, namely grammaticality and faithfulness. Grammaticality necessitates that all grammatical errors should be accurately corrected, while faithfulness requires that the corrections maintain the original textual meaning and syntactic structure. However, the widely-adopted mainstream GEC metrics Bryant et al. (2017); Ye et al. (2023c) indicate the GEC performance by precision, recall, and F scores, which can hardly characterize these critical dimensions of GEC systems, thus hindering the development.
Therefore, we propose CLEME2.0, a more interpretable reference-based evaluation strategy that can describe four fundamental aspects of GEC systems: hit-correction, error-correction, under-correction, and over-correction. The first three aspects are responsible for describing grammaticality, while the last one is for faithfulness since the over-correction edits tend to change the original semantics, especially for LLMs Coyne et al. (2023). To achieve this, CLEME2.0 distinguishes between necessary and unnecessary corrections and disentangles edits into four main types: true positive (TP), necessary false positive (FP), unnecessary false positive (FP), false negative (FN) edits.222True negative edits are not considered in our method. For example in Figure 1, the Hyp.1 makes three necessary edits on the right positions, where [the ] is a TP edit but two of others ([were was] and [for in]) are FP edits since they are not covered in the reference. So Hyp.1 tends to mistakenly correct grammatically errors. On the other hand, Hyp.2 makes extra two FP edits ([ of] and ([century centuries]) since the reference does not correct the right positions, indicating the occurrence of two under-correction phenomena. Additionally, [for for] of the Hyp.2 is considered as an FN edit, which means the occurrence of an under-correction phenomenon. Since the edit disentanglement is based on the chunk partition technique proposed by CLEME, so we dub this strategy as CLEME2.0.
Disentangling edits enables us to investigate concrete dimensions of GEC systems by computing upon an evaluation dataset four disentangled scores: hit-correction, error-correction, under-correction, and over-correction scores. In contrast to mainstream GEC metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a) that reveal the system performance by P/R/F0.5, this disentanglement can provide an interpretable insight into fine-grained dimensions responsible for describing critical characteristics of GEC systems. Then, we integrate these disentangled scores into a comprehensive score using linear weighted summation, placing different emphases on disentangled scores. We leverage the comprehensive score to indicate the system performance from a global perspective. Similar to CLEME Ye et al. (2023c), CLEME2.0 also supports the evaluation based on either correction dependence or correction independence assumptions, providing a flexible option.
Besides, we assume that edits with various extents of modification should affect distinctively the evaluation results. Therefore, we incorporate two edit weighting techniques into CLEME2.0, namely similarity-based weighting Gong et al. (2022) and LLM-based weighting. Specifically, the techniques compute an important weight for each edit using a language model rather than treating each edit equally, thus equipping CLEME2.0 with abilities to capture context semantics and overcome the defect of traditional measures relying on superficial form similarity Kobayashi et al. (2024a).
@@
To verify the effectiveness of CLEME2.0, we conduct extensive experiments on 2 human judgment datasets (GJG15 Grundkiewicz et al. (2015) and SEEDA Kobayashi et al. (2024b)), where our method consistently achieves high correlations. We also demonstrate the robustness of CLEME2.0 by computing the evaluation results based on 6 reference datasets with disparate annotation styles. In summary, our contributions are three folds:
@@
We propose CLEME2.0, a more interpretable evaluation strategy, which is beneficial to reveal crucial characteristics of GEC systems.
We boost CLEME2.0 with two edit weighting techniques, including similarity-based and LLM-based weighting, to overcome the inability of traditional reference-based metrics.
@@ 2 Related Work
Reference-based metrics evaluate GEC systems by referencing manually written materials Ye et al. (2023a, b); Huang et al. (2023); Li et al. (2024a). The M2 scorer Dahlmeier and Ng (2012b) identifies optimal edit sequences between source sentences and system hypotheses, using the F0.5 score. However, this method can inflate scores by manipulating edit boundaries. Bryant et al. (2017) proposed ERRANT, which improves edit extraction with a linguistically-informed alignment algorithm, but it remains language-dependent and biased in multi-reference evaluation. Napoles et al. (2015) introduced GLEU, an n-gram-based metric inspired by BLEU for GEC evaluation. Ye et al. (2023c) proposed CLEME to eliminate bias in multi-reference evaluation by transforming the source, hypothesis, and references into chunk sequences with consistent boundaries, providing unbiased F0.5 scores. Gong et al. (2022) introduce PT-M2, focusing on scoring changed words extracted by the M2 metric.
@@
To overcome the limitations of reference-based metrics, recent research focus on reference-less scoring. Inspired by quality estimation in NMT Liu et al. (2022); Dong et al. (2023), Napoles et al. (2016a) propose Grammaticality-Based Metrics (GBMs) using an existing GEC system or a pre-trained ridge regression model. Asano et al. (2017) enhance GBMs by adding criteria like grammaticality, fluency, and meaning preservation. Yoshimura et al. (2020b) introduce SOME, which uses sub-metrics optimized for manual assessment with regression models. Scribendi Score Islam and Magnani (2021) combines language perplexity and token/Levenshtein distance ratios. IMPARA Maeda et al. (2022) incorporates a Quality Estimator and a Semantic Estimator based on BERT to evaluate GEC output quality and semantic similarity. While reference-less metrics align well with human judgments, they lack interpretability due to the heavy dependence on trained models, thus posing latent risks.
@@ 3.1 Edit Extraction
The first step is edit extraction. Given a source sentence and a target (either hypothesis or reference) sentence , this step is to extract the edits describing the modification from to . Here, we utilize the chunk partition technique from CLEME Ye et al. (2023c) to execute the process of edit extraction. Unlike the traditional metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a), CLEME concurrently aligns all sentences, including the source, the hypothesis, and all the references. This facilitates segmentation of them all into chunk sequences with an equal number of chunks, irrespective of the varying token counts in different sentences, as delineated in Figure 2. It is worth noting that a chunk is a basic edit unit, which can be unchanged, corrected, or dummy (empty) Ye et al. (2023c).
@@ 3.2 Disentangled Scores
For the purpose of computing disentangled scores, we initially disentangle edits into four core types. 1) TP edits refer to the corrected/dummy hypothesis chunks that share the same tokens as the corresponding reference chunks. 2) FP edits are the corrected/dummy hypothesis chunks that have different tokens from those in the corresponding reference chunks wherein the reference chunks are also corrected/dummy ones. 3) FP edits are the corrected hypothesis chunks but their corresponding reference chunks remain unchanged. 4) FN edits indicate the unchanged hypothesis chunks but the corresponding reference chunks are corrected/dummy. It is highlighted that traditional metrics Dahlmeier and Ng (2012a); Bryant et al. (2017); Li et al. (2023b) do not distinguish between FP and FP, treating both as FP, thereby resulting in confusion between error-correction and over-correction. Actually, we have .
@@
Error-correction score.
Conversely, the error-correction score is defined as the ratio of FP edits to all necessary reference edits. This score seeks to evaluate the degree to which systems generate erroneous corrections for grammatical errors. The formula for this score is as follows:
@@
Similarly, the under-correction score is proposed to measure the degree to which systems omit to correct grammatical errors, which is computed as follow:
@@ 3.4 Edit Weighting
Existing reference-based metrics, such as ERRANT and CLEME, depend heavily on superficial literal similarity. This means that, regardless of length or modification, all types of edits have equal weighting in the evaluation scores. This aspect fails to acknowledge that human evaluators might semantically consider the edits’ varying importance levels. Therefore, we introduce two distinct edit weighting techniques to compute the importance weights of edits. These weights are then incorporated into the calculation of the aforementioned disentangled scores as depicted in Equation (1) (4). Take the hit-correction score as a typical example, we reformulate the Equation (1) as follow:
@@
We use PTScore from Gong et al. (2022) to provide edit weights. Through simulating a partially accurate version of the source sentence , PTScore can assign individual weights to edits, in spite of multiple edits in a sentence. Since it performs based on BERTScore Zhang et al. (2019) designed to compute similarity scores for text generation, we call this technique as similarity-based weighting. The computation process is as follows:
where the function is intended to replace a specific chunk of the source with the corrected/dummy hypothesis chunk . Here, denotes the reference sentence. Comprehensive details can be found in Gong et al. (2022).
@@
In light of the impressive semantic understanding capabilities of LLMs Tan et al. (2023); Li et al. (2024b); Yu et al. (2024); Huang et al. (2024); Du et al. (2024); Xu et al. (2024); Qin et al. (2024b), some recent work has sought their use in evaluating assorted NLP tasks Pavlovic and Poesio (2024); Chen et al. (2024); Qin et al. (2024a), GEC evaluations included Sottana et al. (2023). Drawing inspiration from this trend, we employ Llama-2-7B Touvron et al. (2023) as a scorer to acquire the importance score for each edit. These scores range from 1 to 5, with higher scores indicating increased edit importance. We provide the implementation details and the prompting template in Appendix D.
@@ 4.1 Experimental Settings
Reference-based metrics rely on a reference set to establish a system ranking list, the properties of which may significantly influence the performance of the metrics. To investigate the impact of variable reference sets, we assess human consistency across 6 reference datasets. These datasets encompass a range of annotation styles, and a number of human annotators, including CoNLL-2014 Grundkiewicz et al. (2015), BN-10GEC Bryant and Ng (2015) and SN-8GEC Sakaguchi et al. (2016). Notably, SN-8GEC is partitioned into 4 sub-sets, namely E-Minimal, E-Fluency, NE-Minimal, and NE-Fluency. A more thorough breakdown of these datasets and the statistics are provided in Appendix A.

### v2

# CLEME2.0: Towards Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction

@@ Abstract
The paper focuses on the interpretability of Grammatical Error Correction (GEC) evaluation metrics, which received little attention in previous studies. To bridge the gap, we introduce CLEME2.0, a reference-based metric describing four fundamental aspects of GEC systems: hit-correction, wrong-correction, under-correction, and over-correction. They collectively contribute to exposing critical qualities and locating drawbacks of GEC systems. Evaluating systems by combining these aspects also leads to superior human consistency over other reference-based and reference-less metrics. Extensive experiments on two human judgment datasets and six reference datasets demonstrate the effectiveness and robustness of our method, achieving a new state-of-the-art result. Our codes are released at https://github.com/THUKElab/CLEME.
CLEME2.0: Towards Interpretable Evaluation by Disentangling Edits
for Grammatical Error Correction
Jingheng Ye1††thanks: Equal Contribution., Zishan Xu111footnotemark: 1, Yinghui Li1, Linlin Song2, Qingyu Zhou3, Hai-Tao Zheng1,4††thanks: Corresponding Author: Hai-Tao Zheng. (E-mail: zheng.haitao@sz.tsinghua.edu.cn), Ying Shen5, Wenhao Jiang6, Hong-Gee Kim7, Ruitong Liu1, Xin Su8, Zifei Shan8 1Tsinghua University, 2Huazhong University of Science and Technology, 3ByteDance Inc., 4Peng Cheng Laboratory, 5Sun-Yat Sen University, 6Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ), 7Seoul National University, 8Tencent {yejh22,xzs23}@mails.tsinghua.edu.cn
@@ 1 Introduction
The task of Grammatical Error Correction (GEC) automatically detects and corrects grammatical errors in a given text Bryant et al. (2023). A core component of GEC is the development of automatic metrics that can objectively measure model performance Kobayashi et al. (2024b); Ye et al. (2023c). However, proposing appropriate evaluation of GEC has long been a challenging task Madnani et al. (2011), due to the subjectivity Bryant and Ng (2015), complexity Mita et al. (2019) and subtlety Choshen and Abend (2018) of GEC.
Recent research efforts have focused on developing GEC metrics that closely align with human judgements Koyama et al. (2024), whereas the interpretability of these metrics has received less emphasis. We define the interpretability of metrics as their capacity to disclose concerned characteristics of systems, which is crucial for identifying weaknesses in a given GEC system. It is generally recognized that excellent systems should adhere to the principles of grammaticality and faithfulness Bryant et al. (2023). Grammaticality demands that all grammatical errors be accurately corrected, while faithfulness ensures that corrections retain the original meaning and syntactic structure. Nevertheless, the commonly utilized GEC metrics Bryant et al. (2017); Dahlmeier and Ng (2012a) are PRF-based (Precision, Recall, and F scores). We claim that PRF-based metrics fail to effectively capture subtle dimensions of GEC systems, consequently hindering progress. As illustrated in Figure 1, the edits [werewas] and [forin] in Hyp. 1 are regarded as 2 FP + FN edits by ERRANT Bryant et al. (2017) or 2 FP edits by CLEME Ye et al. (2023c). Meanwhile, the edit [of] in Hyp. 2 is categorized as an FP edit for both ERRANT and CLEME. Despite this, these two categories of FP edits carry different implications. The former type is correctly placed but wrongly modified, whereas the latter is incorrectly positioned. The inability to differentiate between these FP edits results in ambiguous interpretations of P/R/F0.5 scores, as they fail to quantify grammaticality and faithfulness.
Thus, we introduce CLEME2.0, an interpretable reference-based metric describing four fundamental aspects of GEC systems: 1) the hit-correction score reflects the degree to which a system accurately corrects grammatical errors, 2) the wrong-correction score denotes the degree of incorrect corrections made, 3) the under-correction score reveals the degree of missing corrections, and 4) the over-correction score measures the degree of excessive corrections. An excellent GEC system should gain a higher hit-correction score and lower wrong-correction, under-correction, and over-correction scores. The initial three aspects assess grammaticality, whereas the over-correction score pertains to faithfulness, given that it often alters the original meaning, a challenge notably observed with LLMs Coyne et al. (2023); Li et al. (2023a). To achieve this, CLEME2.0 first distinguishes between necessary and unnecessary false positive (FP) edits. The idea is that necessary FP edits indicate the system’s wrong-correction degree, while unnecessary FP edits reveal the system’s over-correction degree. As shown in the bottom block of Figure 1, [werewas] and [forin] in Hyp. 1 are regarded as FP edits, while [of] in Hyp. 2 is considered as an FP edit.
As a result, CLEME2.0 establishes a one-to-one relationship between four distinct system aspects and four types of edits: hit-correction v.s. TP, wrong-correction v.s. FP, under-correction v.s. FN, and over-correction v.s. FP. Unlike conventional GEC metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a) that evaluate using P/R/F0.5 scores, it offers a nuanced view into the detailed aspects necessary for characterizing critical features of GEC systems. These separated scores are then consolidated into an overall score via linear weighted summation, giving varying importance to these distinct scores. This aggregate score provides a holistic measure of system performance. Similar to CLEME, our method adopts the chunk partition technique and supports evaluations based on either correction dependence or correction independence assumptions, so we dub the metric as CLEME2.0.
Moreover, we propose that edits of varying modification levels should uniquely influence the evaluation outcomes. For example, corrections involving punctuation are often less significant than corrections of content words. Therefore, we integrate two edit weighting techniques into CLEME2.0, similarity-based weighting Gong et al. (2022) and LLM-based weighting. In particular, these methods compute a specific weight for each edit through a language model rather than assigning equal weight to all edits, thereby enabling CLEME2.0 to grasp contextual semantics and address the limitations of conventional metrics that depend on surface-level form similarity Kobayashi et al. (2024a).
@@
To verify the effectiveness of CLEME2.0, we conduct extensive experiments on two human judgment datasets (GJG15 Grundkiewicz et al. (2015) and SEEDA Kobayashi et al. (2024b)), where our method consistently achieves high correlations. We also demonstrate the robustness of CLEME2.0 by computing the evaluation results based on six reference datasets with disparate annotation styles. In summary, our contributions are three folds:
@@
We introduce CLEME2.0, an interpretable reference-based metric, which is beneficial to reveal crucial aspects of GEC systems.
We enhance CLEME2.0 by incorporating two edit weighting techniques, addressing the limitations of conventional reference-based metrics in capturing semantics.
@@ 2 Related Work
Reference-based metrics evaluate GEC systems by comparing their outputs to manually written references Ye et al. (2022, 2023a, 2023b); Huang et al. (2023); Li et al. (2024c, 2022c, 2022b, d); Ma et al. (2022); Zhang et al. (2023, 2025a); Li et al. (2025c). The M2 scorer Dahlmeier and Ng (2012b) identifies optimal edit sequences between source sentences and system hypotheses, using F0.5 scores. However, this method can inflate scores by manipulating edit boundaries. To mitigate this problem, ERRANT Bryant et al. (2017) improves edit extraction through a linguistically-informed alignment algorithm, but it remains language-dependent and biased in multi-reference evaluation. CLEME Ye et al. (2023c) further provides unbiased F0.5 scores and introduces an extra correction assumption for multi-reference evaluation. PT-M2 Gong et al. (2022) combines PT-based and existing GEC metrics for higher correlations with human judgments.
@@
To overcome the limitations of reference-based metrics, recent studies focus on reference-less scoring. Inspired by quality estimation in NMT Liu et al. (2022); Dong et al. (2023), Napoles et al. (2016a) propose Grammaticality-Based Metrics (GBMs) using an existing GEC system or a pre-trained ridge regression model. Asano et al. (2017) enhance GBMs by adding criteria like grammaticality, fluency, and meaning preservation. Yoshimura et al. (2020) introduce SOME, which uses sub-metrics optimized for manual assessment with regression models. Scribendi Score Islam and Magnani (2021) combines language perplexity and token/Levenshtein distance ratios. IMPARA Maeda et al. (2022) incorporates a Quality Estimator and a Semantic Estimator based on BERT to evaluate GEC output quality and semantic similarity. While reference-less metrics align well with human judgments, they lack interpretability due to the heavy dependence on trained models, thus posing latent risks.
@@ 3.1 Edit Extraction
Given a source sentence and a target (either hypothesis or reference) sentence , we extract the edits describing the modification from to . Here, we utilize the chunk partition technique from CLEME Ye et al. (2023c) to execute the process of edit extraction. Unlike the traditional metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a), CLEME concurrently aligns all sentences, including the source, the hypothesis, and all the references. This facilitates segmentation of them all into chunk sequences with an equal number of chunks, irrespective of the varying token counts in different sentences, as delineated in Figure 2. It is worth noting that a chunk is a basic edit unit, which can be unchanged, corrected, or dummy (empty) Ye et al. (2023c).
@@ 3.2 Disentangled Scores
To compute disentangled scores, we initially disentangle edits into four elementary types. 1) TP edits refer to the corrected/dummy hypothesis chunks that share the same tokens as the corresponding reference chunks. 2) FP edits are the corrected/dummy hypothesis chunks that have different tokens from those in the corresponding reference chunks wherein the reference chunks are also corrected/dummy ones. 3) FP edits are the corrected hypothesis chunks but their corresponding reference chunks remain unchanged. 4) FN edits indicate the unchanged hypothesis chunks but the corresponding reference chunks are corrected/dummy. It is highlighted that traditional metrics Dahlmeier and Ng (2012a); Bryant et al. (2017); Li et al. (2023c) do not distinguish between FP and FP, treating both as FP, thereby resulting in confusion between wrong-correction and over-correction. Actually, we have .
@@
Wrong-correction score.
Conversely, the wrong-correction score is defined as the ratio of FP edits to all necessary reference edits. This score seeks to evaluate the degree to which systems generate erroneous corrections for grammatical errors. The formula for this score is as follows:
@@
Similarly, the under-correction score is proposed to measure the degree to which systems omit to correct grammatical errors, which is computed as follows:
@@
With the disentangled scores indicating disparate aspects of GEC systems, researchers can identify specific weaknesses and implement targeted improvements without expensive human labor.
@@ 3.4 Edit Weighting
Existing reference-based metrics, such as ERRANT and CLEME, depend heavily on superficial literal similarity. This means that, regardless of length or modification, all types of edits have equal weighting in the evaluation scores. This aspect fails to acknowledge that human evaluators might semantically consider the edits’ varying importance levels. Therefore, we introduce two distinct edit weighting techniques to compute the importance weights of edits. These weights are then incorporated into the calculation of the aforementioned disentangled scores as depicted in Equation (1) (4). Take the hit-correction score as a typical example, we reformulate the Equation (1) as follows:
@@
We use PTScore to assign edit weights Gong et al. (2022). Since it performs based on BERTScore Zhang et al. (2019), a tool for evaluating text generation through similarity scores, we refer to this technique as similarity-based weighting. The rationale is to prioritize edits with a more significant modification of the meaning and quality of the text.
By simulating a partially accurate version of the source sentence , PTScore can associate specific weights to edits within a sentence. The computation process is as follows:
where is the reference sentence, while the function is used to replace a specific chunk of the source with the corrected/dummy hypothesis chunk . A positive weight indicates a beneficial correction, whereas a negative value suggests a wrong correction. The absolute value is utilized as the edit weight following (Gong et al., 2022), and the significance111For more detailed analysis, refer to our case study in Section 5.1 and PT-M2 Gong et al. (2022). of an edit in a sentence grows with a larger .
@@
Recent studies have begun investigating the effectiveness of LLM-based evaluation, known for their advanced semantic comprehension Qin et al. (2024b), in assessing various NLP tasks Pavlovic and Poesio (2024); Sottana et al. (2023). Building on this trend, we prompt Llama-2-7B Touvron et al. (2023) to assign edit weights from 1 to 5, where a higher value signifies more critical edits. This methodology is rooted in the idea that LLMs, due to their extensive training on diverse data, are adept at grasping intricate language patterns and text structure. Detailed implementation instructions and the prompting framework are available in Appendix A.
@@ 4.1 Experimental Settings
Reference-based metrics rely on a reference set to establish a system ranking list, the properties of which may significantly influence the performance of the metrics. To investigate the impact of variable reference sets, we assess human consistency across 6 reference datasets. These datasets encompass a range of annotation styles, and a number of human annotators, including CoNLL-2014 Grundkiewicz et al. (2015), BN-10GEC Bryant and Ng (2015) and SN-8GEC Sakaguchi et al. (2016). Notably, SN-8GEC is partitioned into 4 sub-sets, i.e., Expert-Minimal, Expert-Fluency, Non-Expert-Minimal, and Non-Expert-Fluency. A more thorough breakdown of these datasets and the statistics is provided in Appendix B.
@@
Trade-off factors.
We leverage a cross-evaluation search method to identify two optimal sets of trade-off factors for both the corpus and sentence levels. At the corpus level, the factors are assigned as , while for the sentence level, they are adjusted to . The details of the chosen values of trade-off factors can be seen in Appendix B.5.
Evaluation assumptions.
CLEME can evaluate GEC systems based on correction dependence (-dep) or independence (-ind) assumptions. The correction independence assumption offers a more relaxed edit-matching process, implying that systems might yield better scores when multiple references are available. Inspired by this work, CLEME2.0 also supports both assumptions, and we will study their effects on our method.
4.2 Results of GJG15 Ranking
For corpus-level, CLEME2.0-sim-ind achieves the highest average correlations, closely followed by CLEME2.0-sim-dep. CLEME2.0-ind and CLEME2.0-dep can also gain comparable correlations with other metrics, even though they do not utilize any edit weighting techniques. On the other hand, sentence-level metrics exhibit a similar pattern. SentCLEME2.0-sim-dep and SentCLEME2.0-sim-ind achieve the highest Pearson and Spearson correlations, respectively. These results significantly demonstrate the effectiveness and robustness of our proposed method across different settings.
This observation is consistent with recent studies Gong et al. (2022); Ye et al. (2023c). This is because system-level rankings treat each sample equally regardless of edit numbers, mirroring how sentence-level metrics are evaluated. On the other hand, corpus-level metrics emphasize samples with more edits, thus causing the gap between automatic metrics and human evaluation. SentPT-M2 shows superior performance on CoNLL-2014 but performs worse on BN-10GEC, E-Minimal, and NE-Fluency compared to our approach, revealing a lack of robustness of the metric.
Generally, our method aligns more closely with human assessments than existing popular metrics. Notably, our method with similarity-based weighting surpasses unweighted ones, thanks to the integration of semantic factors. However, on E-Minimal and NE-Minimal, weighted and unweighted results are comparable. We suspect this is because these datasets have minimal yet crucial annotations, reducing the possibility of varying weights and the efficacy of edit weighting.
Furthermore, we present comprehensive results of CLEME2.0 on CoNLL-2014 and offer insights into our method for analyzing and identifying weaknesses in GEC systems in Appendix C.5.